AudioCables, Inc., is currently manufacturing an adapter that has a variable cost of $0.50 per unit and a selling price of $1.30 per unit. Fixed costs are $14,000. Current sales volume is 30,000 units. The firm can substantially improve the product quality by adding a new piece of equipment at an additional fixed cost of $6,000. Variable costs would increase to $0.75, but sales volume should jump to 40,000 units due to a higher-quality product. a. What is the current profit and proposed profit of the sales of AudioCables? B. Should AudioCables buy the new equipment? Yes No

Sorry - I figured it out; I unfortunately had my numbers plugged in wrong. Thanks!
Answer #1

Existing equipments

Profit = Revenue - Expense

= Selling Price * Sales Volume - (Fixed Cost + Variable Cost)

= 1.3*30000 - (14000 + 0.5*30000)

= 39000 - 29000

= $10000

New Equipments

Profit = 1.3*40000 - (20000 + 0.75*40000)

= 52000 - 50000

= $2000

No, Audio cable should not buy the new equipment.